If f(2)=1 and f(4)=-2 then what is f(x)?

I just don't know how to do it....

Answer by rothauserc(4718)

we are given two points and need to wrte the equation for the line
the two points are (2,1) and (4,-2)
the standard form for a line is y = mx +b
m = 1 - (-2) / 2 - 4 = -(3/2)
1 = -(3/2)*2 +b
1 = -3 + b
b = 4
the function is f(x) = -(3/2)x +4